### v1

@@ Abstract
Despite the level of attention given to rollups there is limited empirical research about their performance. To address this gap, we conduct a comprehensive data-driven analysis of the late 2023 transaction boom that is attributed to inscriptions: a novel approach to record data onto a blockchain with no outside server needed. Inscriptions were first introduced on the Bitcoin blockchain to allow for the representation of NFTs or ERC-20-like tokens without smart contracts, but were later spread to other blockchains.
This work examines the applications of inscription transactions in Ethereum and its major EVM-compatible rollups and their impact on blockchain scalability during periods of sudden transaction surges. We found that on certain days, inscription-related transactions comprised over 89% on Arbitrum, over 88% on zkSync Era, and over 53% on Ethereum. Furthermore, 99% of these transactions were related to the minting of meme coins, followed by limited trading activity. Unlike L1 blockchains, during periods of transaction surges, zkSync and Arbitrum experienced lower median gas fees, attributable to the compression of L2 transactions for a single L1 batch. Additionally, zkSync Era, a ZK rollup, demonstrated a stronger reduction in fees than optimistic rollups considered in our study: Arbitrum, Base, and Optimism.
keywords:
Inscriptions, Rollup, EVM-compatible blockchains
@@ 1 Introduction
Since the introduction of Bitcoin [23] and later Ethereum [39], blockchains have witnessed increased adoption, predominantly driven by their decentralization principles [5, 27]. This surge has led to the development of a diverse array of decentralized applications (DApps), including decentralized exchanges (DEXes) [12, 3, 18], lending protocols [29, 26], non-fungible tokens (NFTs) for art [13], and other applications such as supply chain management [28, 30], and decentralized governance [18, 3, 20, 21].
@@
In distributed systems, including blockchains, a fundamental trade-off exists where each blockchain can only optimize two out of three essential factors: security, decentralization, and scalability [38]. This dilemma is commonly referred to as the “blockchain trilemma”, a term attributed to V. Buterin [38]. Traditionally, underlying blockchain solutions known as Layer-1 (L1) such as Bitcoin and Ethereum have prioritized decentralization and security at the expense of scalability. To address this challenge, Layer-2 (L2) solutions, built atop an L1, have been proposed. L2 solutions aim to enhance scalability and efficiency while leveraging the L1 blockchain for achieving consensus. Consequently, a variety of Ethereum Virtual Machine (EVM)-compatible chains have emerged, focusing on scalability without compromising decentralization and security. This is facilitated by executing transactions off-chain and subsequently storing results or state transitions on a L1 blockchain (e.g., Ethereum).
However, many of these L2 scaling solutions were not yet fully stress-tested until late 2023, when a spike in transactions occurred: most L2 EVM-compatible chains experienced their all-year heights of transaction volumes in November and December 2023. This sudden transaction spike was caused by a boom in what is known as inscriptions [1]
@@
Inscriptions originated from a new idea that surged in the Bitcoin blockchain - ordinals [37]. Ordinals were first introduced in March 2023 by an anonymous developer known as Domo [6]. In short, this new system allows users to assign unique identifiers to individual satoshis, the smallest unit of Bitcoin, enabling each satoshi to be tracked and traded separately, based on their ordinal number. This concept adds a new dimension to Bitcoin transactions by introducing the ability to treat each satoshi as a non-fungible token (NFT), enabling new applications such as collectible assets, verifiable digital scarcity, and more complex smart contract interactions on the Bitcoin network.
This innovation led to a notable surge in the number of transactions in mid-2023 due to a fear-of-missing-out (FOMO) effect, subsequently causing an increase in transaction fees for all users [36, 37].
@@
Thereafter, the concept of ordinals was extended to other blockchains, such as Ethereum, Arbitrum, Base, Optimism, and zkSync Era, thus leading to the new phenomenon of inscriptions. They enable users to embed arbitrary data directly onto the Ethereum blockchain. This process involves attaching data, such as text, images, or code, to transactions or smart contracts, effectively “inscribing” this information permanently within the blockchain. Users quickly started to use inscriptions to mint and trade tokens, instead of using the already available smart-contract functionality.
Similarly to what happened with Bitcoin ordinals, inscriptions triggered a significant increase in transaction-per-second (TPS) on these chains, reaching yearly transaction highs in December 2023. During this time, inscription transactions represented over $80$% of all network transactions [1].
Due to their apparent popularity, various marketplaces dedicated to trading ordinals and inscriptions began to pop-up [16, 44, 8]. Notably, even the leading crypto-exchange, Binance, supports ordinals and inscriptions [9].
During the inscription boom at the break of 2023 and early 2024, Zero Knowledge (KP) rollups — a type of L2 blockchain — witnessed notably lower gas costs compared to Ethereum, heightening the interest of users, as the cost of participating in the inscription-boom became considerably more affordable when compared to Bitcoin and Ethereum. Nevertheless, some rollups experienced downtime of approximately $78$ minutes [33, 15], leading to users’ concerns about their transactions’ finality.
@@
On March 13, 2024, Ethereum underwent the Dencun upgrade [2], which introduced blobs as temporary data storage. Blobs decreased the gas costs at Ethereum rollups but also raised concerns about the impact on inscription-based tokens ownership since they rely on out-of-the-blockchain applications and therefore presumable assumes data availability [25]. After the 18-day period, blobs disappear from the Ethereum blockchain, as they are designed to support the state verification for rollups. Consequently, inscription-related data might only be available off-chain in the indexer of their platform creator if the rollup chooses not to store all their transaction data.
1.1 Research Questions and Contributions
In this paper we aim to study the performance and user behavior of EVM-compatible chains during the inscription boom. We analyze five chains, namely, Ethereum and its four major rollups—Arbitrum, Base, Optimism, and zkSync Era. To the best of our knowledge, this is the first comprehensive analysis of inscriptions within multiple EVM-compatible chains and the associated user dynamics stemming from the inscription surge observed at the end of 2023 and early 2024.
In particular, we address the following research questions:
[…]
@@ 1.1 Research Questions and Contributions
Our analysis of inscription transactions reveals that the spikes in transaction volume were primarily driven by the minting of new inscription-based meme coins. These tokens were later traded at a fraction of the costs compared to inscriptions on the Bitcoin or Ethereum networks, with prices in the range of \$0.1 to \$1. We found that inscription standards on these chains are based on BRC-20 standards with some modifications to facilitate trading [32, 37, 42] and listing on NFT marketplaces [16, 8, 44].
During multiple daily spikes in transaction counts, even 88% of all transactions were linked to inscriptions, with 99% of these being minting of meme-coins. Specifically, on December 17, 2023, for several hours, over 96% of all transactions on zkSync Era involved the minting of ZRC-20 sync tokens. Other blockchains also exhibited a high percentage of inscription-related transactions. Arbitrum recorded the highest percentage, with over 89%, followed by Ethereum at over 53%, Base at over 37%, and Optimism at the lowest, with over 35%.
[…]
@@
We found that the increased transaction activity resulted in lowered gas fees, positively affecting users of rollup chains. With more L2 transactions compressed into a single transaction batch, the median costs of L2 transactions decreased. Unlike Ethereum, where gas fees increased during periods of transaction surges, median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and zkSync Era. Additionally, zkSync Era, as a ZK rollup, demonstrated a stronger reduction in fees than Arbitrum, an optimistic rollup.
[…]
@@
We observed that across all chains, most on-chain inscription actions were for minting inscriptions rather than selling them. For example, the percentage of inscription minting activities were as follows: Arbitrum at 99.49%, Base at 99.52%, Ethereum at 98.71%, Optimism at 99.73%, and zkSync Era at 99.6%. On zkSync Era, the minting process was completed for three tokens: zrc-20 sync, era-20 bgnt and layer2-20 \$L2, which are traded on zkSync Era for a fraction of the costs compared to the Ethereum and Bitcoin networks. Even with observed prices in the range of \$0.1 to \$1, the costs of minting and listing are covered for the seller. The trading activity for era-20 bgnt token continued after the initial inscription boom.
[…]
We raise concerns related to the applications of blobs for inscriptions. Blobs, temporary data units introduced by Ethereum’s Dencun upgrade [2] are designed to enhance rollup scalability and disappear after 18 days. Rollups are not required to store permanent transaction input data, which inscriptions utilize, leaving users to rely solely on the off-chain indexers run by the inscription creators.
@@
Scientific reproducibility.  Gathering and processing blockchain data can be costly both financially and computationally, even for a single chain. Analyzing multiple chains, as in this study, increases the complexity because most of the data is not even easily accessible. To promote scientific reproducibility and enable other researchers and companies to study blockchain data and inscriptions, we plan to make our datasets and scripts publicly available.
2 Related Work
Novelty of our approach.  To the best of our knowledge, this is the first work to conduct an in-depth analysis of inscriptions in multi EVM-compatible chains and the user dynamics that arose from the inscription boom observed in the end of 2023 and early 2024. While Bitcoin originally set the stage for this user trend, EVM chains have witnessed substantial adoption as well. The absence of analysis in this domain introduces a notable gap in the existing research landscape, which we aim to address in this paper.
3 Background
@@
Scalability challenges and gas price fluctuations in blockchain networks are typically addressed through two primary approaches: Layer-1 (L1) and Layer-2 (L2). Layer-1 is the foundational layer for blockchains, while Layer-2 is a secondary layer that provides scalability and relies on Layer-1 for security. This also introduces rollups, a type of non-custodial Layer-2 blockchain, and provides the necessary background for this paper.
@@
Layer-1 scaling (L1).  This approach involves creating an entirely new blockchain with unique consensus mechanisms and dedicated physical infrastructure to maintain network consensus. Notable L1 blockchains include Bitcoin, Ethereum, Solana, Avalanche, and Polkadot. L1 blockchains typically prioritize decentralization and security, often at the expense of scalability. To address this limitation, L2-rollup solutions were proposed to help scale L1 blockchains.
@@
Layer-2 scaling (L2).  In contrast, L2 scaling adopts a different strategy. It involves executing computations outside the main (L1) blockchain and then recording the results or final state on the underlying chain. Major types of L2 solutions include state (payment) channels, plasma, and rollups. Plasma and state channels focus on moving both data and computation off-chain. Rollups, on the other hand, move computation and state storage off-chain while retaining compressed data for each transaction on the underlying chain.
@@
Rollups.  They function as a form of L2 scaling by conducting computations off-chain and storing the results on the main chain. Essentially, they operate like blockchains, generating blocks and then recording batches of these blocks on the primary chain. However, in the rollup environment, operators are not inherently trusted. This lack of trust means that operators could potentially behave maliciously by halting block production, generating invalid blocks, withholding data, or engaging in other adversarial activities. To mitigate these risks, rollups typically implement robust security measures to ensure the integrity and security of the system.
In that regard, ensuring the correctness of the state in rollups involves two main approaches: optimistic and zero-knowledge proofs (ZKP). Some rollups, such as zkSync Era, employ ZKP, which involve computing cryptographic proofs to validate the correctness of the computations. On the other hand, rollups like Arbitrum and Optimism adopt an optimistic approach, which assumes all statements are true unless someone in the network contests them. Figure 1 illustrates a rollup architecture, highlighting key components such as sequencers and verifiers.
Figure 1: General architecture of rollups—non-custodial L2 blockchain scaling solution.
@@
Sequencers.  These entities are responsible for aggregating transactions and rolling them up to the L1 chain. By bundling transactions, rollups effectively achieve cost savings in terms of gas fees.
@@ 4 Inscriptions and ordinals
Inscriptions and ordinals involve the recording of arbitrary data on the blockchain. On EVM-compatible chains, users encode hex data into the transaction input call data (refer to columns JSON in Table 1), usually also setting the transaction’s from and to attributes with the same user’s addresses. This structure constitutes a self-transfer made from a user to their own address.
@@ 4.1 Operation Types
There are various protocol standards that define the structure of the inscription. BRC-20 was the first standard proposed for Bitcoin, and it includes three main types of operations: deploy, mint, and transfer, each encoded into a single transaction [36]. The protocol standards on EVM-compatible chains extend the set of operations, e.g., claim, list, buy, and sell.
@@
zrc-20 sync transfer $4$’p’: ’zrc-20’, ’op’: ’transfer’, ’tick’: ’sync’, ’amt’: ’4’}
zrc-20 sync mint $4$ data:,{"p":"zrc-20","op":"mint","tick":"sync","amt":"4"}
@@
zrc-20 zkss sell $1000$ data:,{"p":"zrc-20","op":"sell","orderMessage":[{"amt":10000,"nonce":"62...",
"seller":"0x9b...","sign":"0xc5d3...","tick":"zkss","value":0.015}]}
@@
era-20 bgnt list $5$ data:,{"p":"era-20","op":"list","tick":"bgnt","amt":"250","price":"1500000000000000"}
era-20 bgnt buy data:,{"p":"era-20","op":"buy","tx":"0xda..."}
layer-2 \$L2 claim $1000$ data:,{"p":"layer2-20","op":"claim","tick":"\$L2","amt":"1000"}
Deploy.  The deploy action specifies the protocol name, token tick, total supply, and the maximum amount of tokens a user can mint (or claim) per transaction. Table 1 shows examples of inscriptions’ data recorded on-chain. For instance, to deploy an inscription, a transaction containing a deploy action should be recorded on-chain, marking the initiation of an inscription event. In the case of a zrc-20 sync inscription (see first row in Table 1), the protocol specifies a total supply of 21 million inscriptions and a limit of 4 tokens per transaction.
Mint.  After the transaction deploying the inscription is persisted to the chain, users can issue a mint action to actually mint (or in this case, claim ownership) of the tokens. To initiate this, users need to issue a transaction with an input call data encoded to HEX code, specifying the protocol and tick that jointly identify the inscription-based token. In the same transaction, users provide the number of tokens they want to claim (refer to column Limit in Table 1) that must be not higher than the maximum limit specified in deploy operation for the given protocol-tick pair. The off-chain indexer is responsible for assuring the integrity of inscription-based tokens.
Claim.  Similarly to mint, it allows users to mint and claim ownership of a new inscription token. This operation is used, for example, within the layer2-20 standard to mint a new inscription token, \$L2. The \$L2 token is a multi-layer2 inscription token that its inscription creator allows to be minted on multiple blockchains. However, there is just one deploy operation that defines the token’s maximum supply. The off-chain indexer maintained by the creator is responsible for ensuring the integrity of data in these circumstances.

### v2

@@ Abstract
Although rollups have attracted significant attention, there is limited empirical research on their performance under high load. To address this, we present a data-driven analysis of the transaction surge in late 2023 and early 2024, attributed to inscriptions—a method for recording data on the blockchain. Initially introduced on Bitcoin, inscriptions enable the representation of NFTs or ERC-20-like tokens without smart contracts, and have since expanded to other blockchains. This paper examines inscription-related transactions on Ethereum and major EVM-compatible rollups, assessing their impact on scalability during transaction surges. Our results show that, on certain days, inscriptions accounted nearly 90% of transactions on Arbitrum and ZKsync Era, while 53% on Ethereum, with 99% of these inscriptions involving meme coin minting. Furthermore, we show that ZKsync and Arbitrum saw lower median gas fees during these surges. ZKsync Era, a ZK-rollup, showed a greater fee reduction than the optimistic rollups studied—Arbitrum, Base, and Optimism.
@@ 1 Introduction
Since the introduction of Bitcoin [25] and later Ethereum [42], blockchains have seen increased adoption [5, 29]. This growth has led to the development of Decentralized Applicationss (DApps), including Decentralized Exchanges (DEXs) [3, 13, 19], lending protocols [28, 31], Non-Fungible Tokens (NFTs)for art [14], and other applications such as supply chain management [30, 32], and decentralized governance [3, 19, 22, 23].
@@
In distributed systems, including blockchains, a fundamental trade-off has been observed; a blockchain can only optimize two out of the following three essential factors: security, decentralization, and scalability [41]; this is the “blockchain trilemma”, a term attributed to V. Buterin [41]. Traditionally, underlying blockchain solutions known as Layer-1 (L1)such as Bitcoin and Ethereum have prioritized decentralization and security at the expense of scalability. To address this challenge, Layer-2 (L2)solutions, built atop an L1, have been proposed. L2 solutions aim to improve scalability and efficiency while relying on the L1blockchain to heritage security. Consequently, a variety of Ethereum Virtual Machine (EVM)-compatible chains have emerged, focusing on scalability without compromising decentralization and security. This is facilitated by executing transactions off-chain and subsequently storing results or state transitions on an L1blockchain.
@@
However, many of these L2scaling solutions were not yet fully stress-tested until late 2023, when a spike in transactions occurred: most L2 EVM-compatible chains experienced their all-year heights of transaction volumes in November and December 2023. This sudden transaction spike was caused by a boom in what is known as inscriptions [1]. Inscriptions originated from a new idea that surged in the Bitcoin blockchain — ordinals [40]. Ordinals were first introduced in March 2023 by an anonymous developer known as Domo [6]. In short, this new system allows users to assign unique identifiers to individual satoshis, the smallest unit of Bitcoin, enabling each satoshi to be tracked and traded separately, based on their ordinal number. This concept adds a new dimension to Bitcoin transactions by introducing the ability to treat each satoshi as a NFT, enabling new applications such as collectible assets, verifiable digital scarcity, and more complex smart contract interactions on the Bitcoin network. This led to a notable surge in the number of transactions in mid-2023 due to a Fear-Of-Missing-Out (FOMO)effect, which subsequently caused an increase in transaction fees for all users [39, 40].
@@
Subsequently, the concept of ordinals was extended to other blockchains, such as Ethereum, Arbitrum, Base, Optimism, and ZKsync Era, thus leading to the new phenomenon of inscriptions. They enable users to embed arbitrary data directly onto these blockchains. This process involves attaching data, such as text, images, or code, to transactions or smart contracts, effectively “inscribing” this information permanently into the blockchain. Users quickly started to use inscriptions to mint and trade tokens, instead of using the already available smart contract functionality. Similarly to what happened with Bitcoin ordinals, inscriptions triggered a significant increase in Transaction-per-Second (TPS)on these chains, reaching yearly transaction highs in December 2023. During this time, inscription transactions represented over $80$% of all network transactions [1]. Due to their apparent popularity, various marketplaces dedicated to the trading of ordinals and inscriptions began to pop-up [8, 17, 48]. In particular, even the leading crypto-exchange, Binance, supports ordinals and inscriptions [9].
@@
During the inscription boom at the break of 2023 and early 2024, Zero Knowledge (ZK)rollups — a type of L2blockchain — witnessed notably lower gas costs compared to Ethereum, increasing the interest of users, as the cost of participating in the inscription boom became considerably more affordable compared to Bitcoin and Ethereum. However, some rollups experienced downtime of approximately $78$ minutes [16, 36]. On March 13, 2024, Ethereum underwent the Dencun upgrade [2], which introduced blobs as temporary data storage. Blobs decreased gas costs for Ethereum rollups, but also raised concerns about the impact on inscription-based tokens ownership, since they rely on out-of-the-blockchain applications and therefore presumably assume data availability [27]. Note that after the 18-day period, blobs may disappear from the Ethereum blockchain, as they are designed to support state verification for rollups. Consequently, inscription-related data might only be available off-chain in the indexer of their platform creator if the rollup chooses not to store all their transaction data.
We summarize the contributions of this paper as follows.
@@
Causes of transaction surge at the end of 2023.  Our analysis reveals that the surge in transaction volume were primarily driven by the minting of new inscription-based meme coins. These tokens were later traded for a fraction of the costs compared to inscriptions on the Bitcoin or Ethereum networks, with prices in the range of then cents to one US dollar. We found that the inscription standards on these chains are based on the BRC-20 standards with some modifications to facilitate trading [35, 40, 46] and listing on NFTmarketplaces [8, 17, 48]. Specifically, on December 17, 2023, for several hours, over 96% of all transactions on ZKsync Era involved the minting of ZRC-20 sync tokens. Other blockchains also exhibited a high percentage of inscription-related transactions. Arbitrum recorded the highest percentage, with over 89%, followed by Ethereum at over 53%, Base at over 37%, and Optimism at the lowest, with over 35%.
@@
Impact on EVMblockchain performance and cost efficiency.  Given that certain L2 solutions experienced downtime, and a significant portion of transactions during late 2023 and early 2024 were inscriptions, we seek to evaluate the performance of these L2 scaling solutions during this stress-test. We found that the increased transaction activity resulted in a decrease in gas fees, positively affecting the users of the rollup chains. With more L2 transactions compressed into a single transaction batch, the median costs of L2 transactions decreased. Unlike Ethereum, where gas fees increased during periods of transaction surges, median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and ZKsync Era. Additionally, ZKsync Era, as a ZK rollup, demonstrated a stronger reduction in fees than Arbitrum, an optimistic rollup.
@@
Speculative purchases.  We observed that across all chains, most on-chain inscription actions were for minting inscriptions rather than selling them. For example, the percentage of inscription minting activities were as follows: Arbitrum at 99.49%, Base at 99.52%, Ethereum at 98.71%, Optimism at 99.73%, and ZKsync Era at 99.6%. On ZKsync Era, the minting process was completed for three tokens: zrc-20 sync, era-20 bgnt and layer2-20 \$L2, which are traded on ZKsync Era for a fraction of the costs compared to the Ethereum and Bitcoin networks. Even with observed prices in the range of \$0.1 to \$1, the costs of minting and listing are covered for the seller. The trading activity for the era-20 bgnt token continued after the initial inscription boom.
@@
Scientific reproducibility.  Gathering and processing blockchain data can be costly both financially and computationally, even for a single chain. Analyzing multiple chains, as in this study, increases complexity because most of the data are not even easily accessible. To promote scientific reproducibility and enable other researchers and companies to study blockchain data and inscriptions, we plan to make our datasets and scripts publicly available.
2 Background on Rollups
@@
Scalability challenges and gas price fluctuations in blockchain networks are typically addressed through Layer-2 (L2)solution. While, L1is the foundational layer for blockchains, L2is a secondary layer that provides scalability and relies on L1for security. This also introduces rollups, a type of non-custodial L2blockchain, and provides the necessary background for this paper.
@@
Layer-1 (L1).  This approach involves creating an entirely new blockchain with unique consensus mechanisms and dedicated physical infrastructure to maintain network consensus. Notable L1blockchains include Bitcoin, Ethereum, Solana, Avalanche, and Polkadot. L1blockchains typically prioritize decentralization and security, often at the expense of scalability. To address this limitation, L2-rollup solutions were proposed to help scale L1blockchains.
@@
Layer-2 (L2).  In contrast, L2scaling adopts a different strategy. It involves executing computations outside the main (L1) blockchain and then recording the results or final state on the underlying chain. Major types of L2solutions include state (payment) channels, plasma, and rollups. Plasma and state channels focus on moving both data and computation off-chain. Rollups, on the other hand, move computation and state storage off-chain while retaining compressed data for each transaction on the underlying chain.
@@
Rollups.  They function as a form of L2scaling by performing computations off-chain and storing the results on the L1chain. They operate like blockchains, generating blocks and then recording batches of these blocks on the L1chain. However, in the rollup environment, operators are not inherently trusted. This lack of trust means that operators could potentially behave maliciously by halting block production, generating invalid blocks, withholding data, or engaging in other adversarial activities [10]. To mitigate these risks, rollups typically implement robust security measures to ensure the integrity and security of the system. In that regard, ensuring the correctness of the state in rollups involves two main approaches: optimistic and Zero Knowledge Proof (ZKP). Some rollups, such as ZKsync Era, employ ZKP, which involve computing cryptographic proofs to validate the correctness of the computations. On the other hand, rollups like Arbitrum and Optimism adopt an optimistic approach, which assumes all statements are true unless someone in the network contests them.
@@
Sequencers.  These entities are responsible for aggregating, ordering, and rolling up the transactions to the L1chain. By bundling transactions, rollups effectively achieve cost savings in terms of gas fees.
@@ 3 Inscriptions and Ordinals
Inscriptions and ordinals involve the recording of arbitrary data on the blockchain. On EVM-compatible chains, users encode HEX data into the transaction input call data (refer to column JSON in Table 1), usually also setting the transaction’s from and to attributes with the same user addresses. This structure constitutes a self-transfer made from a user to their own address.
@@ 3.1 Operation Types
There are various protocol standards that define the structure inscriptions. BRC-20 was the first standard proposed for Bitcoin, and includes three main types of operations: deploy, mint, and transfer, each encoded in a single transaction [39]. The protocol standards on EVM-compatible chains extend this set of operations by adding, for example, claim, list, buy, sell, among others.
@@
zrc-20 sync transfer $4$"p": "zrc-20", "op": "transfer", "tick": "sync", "amt": "4"}
zrc-20 sync mint $4$"p":"zrc-20","op":"mint","tick":"sync","amt":"4"}
@@
zrc-20 zkss sell $1000$"p":"zrc-20","op":"sell","orderMessage":[{"amt":10000,"nonce":"62...",
@@
era-20 bgnt list $5$"p":"era-20","op":"list","tick":"bgnt","amt":"250","price":"1500000000000000"}
era-20 bgnt buy {"p":"era-20","op":"buy","tx":"0xda..."}
layer-2 \$L2 claim $1000$"p":"layer2-20","op":"claim","tick":"\$L2","amt":"1000"}
Deploy.  The deploy action specifies the protocol name, token tick, total supply, and the maximum amount of tokens a user can mint (or claim) per transaction. Table 1 shows examples of inscriptions’ data recorded on-chain. For example, to deploy an inscription, a transaction containing a deploy action should be recorded on-chain, marking the initiation of an inscription event. In the case of a zrc-20 sync inscription (see first row in Table 1), the protocol specifies a total supply of 21 million inscriptions and a limit of 4 tokens minted per transaction.
Mint.  After the transaction that deploys the inscription persists in the chain, users can issue a mint action to actually mint (or in this case claim ownership) of the tokens. To initiate this, users need to issue a transaction with an input call data encoded to HEX code, specifying the protocol and tick that jointly identify the inscription-based token. In the same transaction, users provide the number of tokens they want to claim (refer to column Limit in Table 1) that must not exceed the maximum limit specified in deploy operation for the given protocol-tick pair. The off-chain indexer is responsible for ensuring the integrity and user balances of inscription-based tokens.
Claim.  Similarly to mint, it allows users to mint and claim ownership of a new inscription token. This operation is used, for example, within the layer2-20 standard to mint a new inscription token, \$L2. The \$L2 token is a multi-L2inscription token that its inscription creator allows to be minted on multiple blockchains. However, there is just one deploy operation that defines the token’s maximum supply. The off-chain indexer maintained by the creator is responsible for ensuring the integrity of data in these circumstances.